As soon as traffic lights changed green, Abdullah accelerates uniformly with its motorcycle at 2.6m/s² for 5 seconds. He then maintain a steady velocity of for ⅔ of 150m journey. How far does he travel from the traffic light junction until he reached a maximum velocity?

Grade:12th pass

2 Answers

Saurabh Koranglekar
4 years ago
Dear student

He traveled for 5 sec with acceleration so at 5 seconds he reached max velocity

S = 0.5*a*t*t
= 1.3*25
= 32.5 meters

Regards
Khimraj
4 years ago
S = 0.5*a*t*t= 1.3*25= 32.5 meters
…...................................................................................